OpenStudy (anonymous):

Convert the rate 5 ft per 3 seconds to yards per hour

OpenStudy (anonymous):

the big thing with conversions is to watch your units, i'm sure you understand the ratio's of like 60 seconds in 1 minute. the idea is to arrange your ratio of 60 sec to 1 min so one unit will cancel out with another like this... 

OpenStudy (anonymous):

\[\frac{ 5 feet }{ 3 seconds} \times \frac{ yard }{ 3 feet}\times \frac{ 60 seconds }{ 1 minute}\times \frac{ 60 \min }{ 1 hr}\]